1 L of air weights 1 .293 g at 0 degree c and 1 atm pressure . at what temperature 1 L of air at 1 atm pressure will weight 1 g

Grade:11

1 Answers

Arun
5 years ago
Applying Ideal gas equation in both case and equating them :-P1V1/ P2V2 = n1RT1/n2RT2Canceling P,V and R since these are not changing we get,n1T1 = n2T21.293/M * 273 = 1/M *TM gets cancelledT = 352.989 K
Regards
Arun